OpenStudy (moonlitfate):

Find the derivative of the function. y= 1/2 [ x ((81-x^2)^(1/2)) + 81 arcsin (x/9)]

OpenStudy (moonlitfate):

The first one.

OpenStudy (moonlitfate):

\[y = \frac{ 1 }{ 2 } [ x \sqrt{81-x^2} + 81 \arcsin \frac{ x }{ 9 }\]

OpenStudy (aum):

Apply product rule and chain rule. What is the derivative of x? What is the derivative of \(\sqrt{81-x^2} \)? What is the derivative of \(\arcsin \frac{ x }{ 9 } \)?

OpenStudy (moonlitfate):

x' = 1 derivative of \[\arcsin \frac{ x }{ 9 } = \frac{ 9 }{ \sqrt{1 - \frac{ x^2 }{ 81 }} }\]

OpenStudy (moonlitfate):

@aum - The part with the chain rule is messing me up. :/

OpenStudy (moonlitfate):

I know \[\sqrt{81-x^2} = (81-x^2)^{1/2}\]

OpenStudy (aum):

Derivative of (something)^n = n * (something)^(n-1) * derivative of (something) \[ \frac {d}{dx}\sqrt{81-x^2} = \frac {d}{dx}(81-x^2)^{1/2} = \frac 12~*~(81-x^2)^{-1/2}~*~\frac {d}{dx}(81-x^2) = ? \]

OpenStudy (moonlitfate):

\[\sqrt{\frac{ 2 }{ 81-x^2 }}\]

OpenStudy (moonlitfate):

Forgot to multiply it by 2x...

OpenStudy (aum):

The derivative in the last part of my previous reply is: (-2x). Multiply all three factors.

OpenStudy (moonlitfate):

Oh, right. I had -2x on the paper I'm trying to work this out on.

OpenStudy (aum):

\[ \frac {d}{dx}\sqrt{81-x^2} = \frac {-x}{\sqrt{81-x^2}} \]

OpenStudy (aum):

Similarly, derivative of arcsin(something) = 1/ sqrt(1 - (something)^2) * derivative of (something) \[ \frac {d}{dx}\arcsin(x/9) = \frac {1}{\sqrt{1 - (x/9)^2}} ~*~ \frac {d}{dx}(x/9) = \\ \frac {1}{\sqrt{1 - (x/9)^2}} ~*~ \frac 19 = \frac{9}{\sqrt{81-x^2}} * \frac 19 ~=~ \frac{1}{\sqrt{81-x^2}} \]

OpenStudy (aum):

Now you can put them all together and find y' using the product rule.

OpenStudy (anonymous):

hello again you got this ?